OpenStudy (anonymous):

What is the limit as x goes to zero for: ((sin(2x))/(7x)?

OpenStudy (anonymous):

pls explain any answer - thanks a ton!@

OpenStudy (cruffo):

Remember the special limit for sin(x)/x ?

OpenStudy (anonymous):

=1

OpenStudy (anonymous):

The easiest way to do this problem is to use L'Hopital's rule. If you plug in 0 into the fraction you get 0/0, which is the parameter for using L'Hopital's. So the rule goes, if top and bottom equal zero, you can take the derivative of top and bottom until one does not go to zero. \[d(\sin(2x))=2\cos2x\] \[d(7x)=7\] so you get \[2\cos(2x)/7\] where x goes to zero, so plug it in and you get 2/7. The limit is 2/7

OpenStudy (cruffo):

Good. It works the same for sin(nx)/nx were n is some number.

OpenStudy (anonymous):

cool, hope my explanation wasn't too complicated

OpenStudy (anonymous):

one sec< while I process - thanks all> I have abt 5 of these q's and I really want to get them

OpenStudy (cruffo):

Have you used L'Hopitals Rule yet?

OpenStudy (anonymous):

True, if you haven't dealt with L'Hopital's yet then you should go with cruffo's explanation

OpenStudy (anonymous):

yes, i think so

OpenStudy (anonymous):

how does the derivative of sin(2x) become 2cos(2x)?

OpenStudy (cruffo):

This problem is a typical "first time using limits" problem, in which case you would go with an algebraic solution, since you haven't covered derivative rule yet.

OpenStudy (anonymous):

that has really confused me :-) sorry

OpenStudy (anonymous):

we just covered the pinching theorem

OpenStudy (cruffo):

Ok, put PJW_Largo's solution aside for about 3 weeks :)

OpenStudy (anonymous):

yeah sorry about that you'll probably get to those derivatives soon though

OpenStudy (anonymous):

no worries; we pull out the number next to the x ( not by the trig function)

OpenStudy (cruffo):

Try this, \[\frac{sin(2x)}{7x} \] \[= \frac{1}{7} \frac{\sin(2x)}{x}\] \[ = \frac{1}{7}\cdot\frac{2}{2}\cdot\frac{\sin(2x)}{x}\]

OpenStudy (anonymous):

forget l'hopital because you have not got there yet. you know \[lim_{x->0}\frac{sin(x)}{x}=1\] so \[lim_{x->0} \frac{sin(2x)}{7x}=\frac{2}{7}\]

OpenStudy (anonymous):

yay! I get it!!!!

OpenStudy (anonymous):

rock n roll!!!

OpenStudy (anonymous):

ok, 5 more here we go!

OpenStudy (cruffo):

:)